OpenStudy (anonymous):

The solution for a system of two linear equations is (0, 0). Which equation below could be one of the equations in the system?

OpenStudy (anonymous):

y= 1/3x + 4 y = –3x y = 4x – 5 y = 5 – x

OpenStudy (anonymous):

Plug in 0 for x and 0 for y in each answer choice. Only one of those answers will be the same on both sides of the = sign after plugging in and simplifiying

OpenStudy (anonymous):

I tried to but I couldn't get an answer for any of them...

OpenStudy (anonymous):

Plug in 0's for x an y in y= 1/3x + 4 What do you get? 0 = 1/3*0 + 4 0 = 4 incorrect..so it's not A keep going

OpenStudy (anonymous):

y = -3 which is 0 = -3(0). so, 0 = 0? is that it?

OpenStudy (anonymous):

that's it!

OpenStudy (anonymous):

Awesome! thanks!